If 2^x = 4^y = 8^z and (1/2x)+(1/4y)+(1/8z)=4, find the value of x.
Answer by Alan3354(69443)

If 2^x = 4^y = 8^z and (1/2x)+(1/4y)+(1/8z)=4, find the value of x.
-------------------------
2^x = 2^(2y) = 2^(3z)
x = 2y = 3z
y = x/2
z = x/3
====================
(1/2x)+(1/4y)+(1/8z)=4
Is it (1/2)x or 1/(2x) ?
-------------
1/(2x) + 1/(4*(x/2)) + 1/(8*(x/3)) = 4
1/(2x) + 1/(2x) + 1/(8x/3) = 4
1/2 + 1/2 + 3/8 = 4x
4x = 11/8
x = 11/32